### v2

# DNA and Human Language: Epigenetic Memory and Redundancy in Linear Sequence

Abstract
DNA is often described as the “language of life”, but whether it possesses formal linguistic properties remains unresolved. Here, we present the first empirical evidence that DNA sequences exhibit core linguistic features—specifically, functional and information redundancy—through comprehensive analysis of genomic and epigenetic datasets. By mapping DNA sequences into a linguistic feature space, we demonstrate that fixed-length (41 bp) DNA segments encode information analogously to human language, with redundancy contributing to signal stability in aqueous intracellular environments. Moreover, we provide the first evidence of one-dimensional epigenetic memory, showing that linear DNA sequences can maintain epigenetic marks such as 6mA methylation—contrasting with models focusing on epigenetic memory transmission via 3D chromatin organization[1]. Our tailored linguistic mapping strategy also addresses persistent challenges in genomic data processing, significantly improving data cleaning and feature extraction. Together, these findings establish a conceptual paradigm that bridges molecular information encoding and linguistic theory, laying the foundation for next-generation large language models (LLMs) specifically tailored to DNA—marking a shift at the interface of molecular biology, information theory, and artificial intelligence (AI).
*Email: db.wang@njau.edu.cn (D. Wang)
1 Introduction
Early studies demonstrated that DNA-binding enzymes primarily recognize specific DNA sequences through interactions within the major groove of the double helix[2]. This mechanism can be conceptually compared to the encoding and decoding of information in human language, where DNA-binding enzymes act as “readers” that recognize and interpret specific instructions embedded within DNA sequences.While this analogy enhances our understanding of molecular recognition, it also raises a compelling question: Do DNA and human languages (such as Chinese or English) exhibit parallels in their informational structure or coding logic? More intriguingly, could a formal connection be established between these two systems of information encoding?
For many years, Scientists have dreamed of linking the ‘language of life’—DNA—with human languages to deepen our understanding of DNA’s encoding and transmission mechanisms[3]. In 1994, Eugene Stanley observed that the frequency distribution of symbols in DNA sequences conforms to Zipf’s law[4], a statistical feature also found in natural language. This similarity sparked an interest in exploring the relationship between DNA and language. If we view organisms as carriers of language, their various components might correspond to words, phrases, and sentences in linguistic structures. This analogy suggests that natural language processing (NLP) methods could be applied to decode DNA. From an information theory perspective, redundancy enhances the accuracy of information transmission. Redundant information refers to "extra information added to the original data to ensure the correct reception of a message even in the presence of noise and other disturbances [5]." Redundancy in communication means that hearing one word helps us predict the next, ensuring that occasional missed or misheard words do not impede understanding. Redundant languages are more robust and adaptable. This raises the question: does short fixed-length DNA, as the ‘language of life’, also exhibit redundancy? While this remains unresolved, redundancy phenomena—such as the ‘many-to-one’ relationship between codons and amino acids, multiple gene copies, and genetic compensation mechanisms [6, 7]—contribute to the stability of life processes. With rapid advances in artificial intelligence, deep learning models can now automatically identify key information based on linguistic features, contextual information, and biological context—despite redundancy. Researchers are increasingly applying natural language algorithms to analyze DNA sequence information. For instance, when predicting DNA methylation on N6-adenine (6mA) methylation sites, algorithms such as LSTM, BiLSTM, and Transformer have significantly improved accuracy[8, 9, 10]. However, in some cases, model performance remains limited (e.g. below 90%)[11], often due to dataset quality and preprocessing issues. Unlike natural languages, DNA consists of only four symbols, which poses unique challenges in data cleaning—a pressing issue in bioinformatics that remains unresolved.
To understand the information storage mechanisms of DNA sequences and overcome the challenges associated with cleaning DNA sequence datasets, we address these issues simultaneously. If DNA datasets could be manually cleaned in the same way as Chinese or English sentences, this would suggest that the information storage mechanism of DNA could be deciphered by humans.In this study, we take 6mA methylation as an entry point to investigate the information-encoding logic embedded in DNA sequences. Using the BERT pre-trained model as a tool, we construct a custom dictionary to map DNA into the feature space of the Chinese language, thereby simulating the “transition state” of bases during the methylation process. Our results demonstrate that using Chinese to simulate the base “transition state” in DNA methylation reaction is feasible. The results indicate that short DNA segments contain functionally meaningful signal encodings with redundant features. This finding suggests that DNA harbors a symbolic information structure analogous to human language—capable of supporting memory maintenance, signal stability, and error self-correction within biological systems. It not only provides theoretical support for transcending traditional biochemical perspectives, but also lays the groundwork for constructing DNA-specific large language models that integrate symbolic encoding with molecular dynamics.
2 Language feature representation of DNA sequences
In DNA symbol representation, the nucleotides A, T, C, and G at different positions are treated as the same. However, we argue that the representation of nucleotide symbols at different positions during the reaction should not be identical. Therefore, we constructed a dictionary that maps the DNA symbols A, T, C, and G to Chinese characters to simulate the "transition state" of the bases during the methylation reaction. This mapping is achieved through a predefined dictionary that facilitates the transformation of DNA into human language (Chinese). We believe that different representations of A, T, C, and G at different positions can better simulate the nucleotides complex states during the methylation process. Although we do not know the exact detailed states, we assume that they differ from the normal state.
To verify the effects of mapping DNA language to the Chinese language feature space, we selected 6mA methylation data from Arabidopsis thaliana (A. thaliana) and Drosophila melanogaster (D. melanogaster), and Xanthomonas oryzae pv. Oryzicola BLS256 (Xoc BLS256) to represent DNA language. According to previous studies, a sequence length of 41bp has been found to achieve satisfactory prediction performance[12], indicating that a sequence length of 41bp is sufficient to capture the key nucleotide information for 6mA. Therefore we chose a sample length of 41bp. To avoid errors caused by single-molecule real-time (SMRT) sequencing technology and the effects of many duplicate samples, we performed filtering and deduplication to generate positive samples. Negative samples were randomly generated from the genome in equal amounts, with each negative sample differing from the known methylation site by at least 20 nucleotides. Finally, we obtained the A. thaliana dataset 1.0, the D. melanogaster dataset 1.0, and the Xoc BLS256 dataset 1.0.
We divided each species’ dataset 1.0 into a training set (80%) and a testing set (20%), ensuring no overlap between them. We performed 5-fold cross-validation to avoid errors in model training. The model evaluation metrics included six indicators: ROC curve, Area Under the Curve (AUC), accuracy(ACC), precision(P), recall(R), and F1 score(F1). Ultimately, by mapping DNA features to the linguistic space with an agreed-upon dictionary, we observed that the BERT model demonstrated superior discrimination performance for 6mA, achieving an AUC of over 0.90, with the value for Drosophila close to 0.97 (Fig.1a). By comparing the 4-feature description with the 164-feature description after dictionary mapping, both show similar ROC curves and AUC (Fig. 1a). There is no significant difference in ACC, P, R, or F1 between the A. thaliana and D. melanogaster datasets (Fig. 1b). On Xoc BLS256, the 164-feature inputs significantly improved ACC, F1, and P (Fig. 1b). Mapping DNA language to Chinese via a dictionary did not reduce the key information encoded in DNA sequences; rather, it tended to highlight critical nucleotide information in Xoc BLS256.
Fig. 1: Prediction of DNA 6mA loci represented by different linguistic features. a, ROC curves for 5 independent experiments comparing 4-feature and 164-feature models, along with the AUC. b, Comparison of 4-feature and 164-feature models in terms of ACC, F1, P, R, and MCC. c, Comparison of 164-feature models using different dictionaries, focusing on ACC, F1, P, R, and MCC. d, A 2D representation of input features derived from the BERT pretrained model.
3 Different linguistic features do not affect key information in a DNA sequence
To further explore the impact of different linguistic mapping features on the 6mA sequences, we evaluated the effect of feature vector representations from different mapping dictionaries on DNA methylation feature learning. To investigate how different linguistic features impact the 6mA prediction model, we utilized three dictionaries: Dictionary 1 (Dict1), Dictionary 2 (Dict2), and Dictionary 3 (Dict3), each containing 164 unique Chinese characters. The Chinese characters in these three dictionaries do not overlap, and their visualization is shown in Fig. 1d. As shown in Fig. 1c, we observed that the use of different dictionaries did not impact the critical information in the DNA sequence, and there were no significant differences in ACC, F1, P, R, or MCC.
4 6mA methylation Information Extraction and Storage Methods in DNA
Fig. 2: Methylation information storage, represented by motifs ( (P(pos)$\geq 90\%$), within a single sequence in A. thaliana, D. melanogaster, and Xoc BLS256. (P(pos) represents the proportion of the motif in positive samples.)
The above experiments showed that mapping with the dictionary had no impact on the information regarding 6mA methylation. However, the new feature representation facilitates the use of rule mining algorithms to extract nucleotide information related to 6mA methylation, which is equivalent to identifying motifs. A DNA motif refers to short, similar repeating patterns of nucleotides with biological significance[13]. Mapping nucleotides to a linguistic space frees us from the limitation of one-dimensional sequence information, bringing it closer to the true 3D state of nucleotides in cells. DNA was converted into a Chinese character representation (Dict1), and the Rule Algorithm was then applied to identify rules (motifs). These rules indicate that when certain key nucleotides appear, the central adenine(A) will be methylated.
A methylated DNA sequence is expected to have its methylation features (motifs) appear more frequently in positive samples and less frequently in negative samples. We calculated the proportion of these motifs in both positive and negative samples, reflecting the tendency of motifs to appear in different categories, as recorded in Supplementary Tables 3-5. At the same time, these methylation features serve as discriminative features for deep learning algorithms to distinguish between positive and negative samples. Interestingly, even without prior knowledge of motifs, GGAGG (methylation site underlined) and CCAN(9)TGCT mined from Xoc BlS256 were consistent with those recorded in the REBASE[14] database (Supplementary Fig. 2 and Supplementary Table 5).
The high proportion of motifs in positive samples often highlights their significance, which led us to analyze the distribution of motifs within individual sequences. As shown in Fig. 2, we present the storage pattern of motifs (P(pos)$\geq 90\%$) within a single sequence. We were surprised to find that a single sequence (equivalent to a sentence) could match multiple motifs (Fig. 2), suggesting that organisms reserve multiple redundant information fragments within fixed-length DNA sequences to enhance the stability of information transmission. Moreover, the efficacy of each redundant message in conveying information exhibits subtle differences. For example, in A. thaliana, the motif A..G..CG proves more effective than A…..A..G..C (Fig. 2a). Such redundancy in DNA storage mechanisms is reminiscent of redundancy in human language. DNA and human language share many similarities: i) Both DNA and human language utilize redundancy to achieve error tolerance, preventing information transmission failure. ii) Both DNA and human language allow the same information to be expressed in different forms, enhancing flexibility. iii) Redundancy provides both DNA and human language with a certain degree of openness, enabling the emergence of new functions or interpretations in the future. A detailed comparison of redundancy functions in DNA and human language can be found in Supplementary Table 27.This strategy of preserving multiple motifs recognized by methyltransferases provides organisms with a significant advantage in maintaining signal stability, even as DNA sequences undergo random and dynamic fluctuations in a liquid medium. In Supplementary Data 1–3, we provide additional records of information storage in single sequences.
5 Cleaning the Dataset to Verify 6mA Methylation Information
To assess the critical motifs, thresholds of 0.95, 0.90, 0.85, 0.80, and 0.75 were set based on the proportion of these motifs in positive samples (Supplementary Tables 3-5). After deleting negative samples containing these motifs, we compared the results with an equally sized group of randomly deleted negative samples. Since the positive samples were measured by SMRT, we were unable to determine which ones might be false positives. Therefore, only the negative samples were considered for testing.
We observed a significant improvement in the model’s performance when negative samples were removed by motifs, whereas randomly removing negative samples under the same conditions did not show similar improvements (Fig. 3). After excluding false negative samples, the AUC of the models for the three species significantly increased. At a threshold of 0.75, the model’s AUC reached 0.99, 1.00, and 0.99 for A. thaliana, D. melanogaster, and Xoc BlS256, respectively (Fig. 3a). The model’s ACC, F1, P, and R demonstrated improvements compared to the random deletion of an equal number of false-negative samples (Fig. 3b). These results demonstrate that the motifs are indeed the key information for the positive samples. Furthermore, after randomly deleting negative samples, as the imbalance between positive and negative samples increased, the accuracy (ACC), F1 score, precision (P), and recall (R) were also affected (Fig. 3b)
Fig. 3: Assessment of critical 6mA methylation information by deleting negative samples in the training data using different thresholds and by random deletion. a, ROC curves for 5 independent experiments using 164-feature and their average AUC. b, Comparison of 164 features in terms of Acc, F1, P, and R. c, Changes in positive and negative sample counts when different thresholds are used to delete negative samples.
6 Interpretation of cross-validation results
Cross-validation between species reflects the differences in 6mA methylation information across species. We predicted the set-aside data using the model trained on each species’ dataset 1.0. These set-aside data was positive samples from each species. As shown in Fig. 4a, we found that the prediction accuracies of the models for each species on their set-aside data were 0.847 for A. thaliana, 0.939 for D. melanogaster, and 0.917 for Xoc BlS256, respectively. The model trained on the A. thaliana dataset 1.0 achieved a much higher recognition accuracy for D. melanogaster (0.952) than for A. thaliana (0.847), while the model trained on the D. melanogaster dataset 1.0 performed worse for A. thaliana (0.663) than for D. melanogaster (0.939). In Supplementary Fig.2 and Supplementary Tables 3-4, it can be found that many motifs in A. thaliana tended to be a subset of D. melanogaster motifs, which accounts for the observed performance differences. The motifs of Xoc BLS256 differed significantly from those of both D. melanogaster and A. thaliana (as illustrated in Supplementary Fig. 2 and Supplementary Tables 3-5). As a result, the Xoc BLS256 recognition model achieved only an accuracy of 0.4 and 0.679 for A. thaliana and D. melanogaster on the set-aside data respectively.
Fig. 4: Assessment of critical 6mA methylation information by deleting negative samples in the training data using different thresholds and by random deletion. a. Cross-species ACC performance validation: Using Dataset 1.0 and Dictionary 1 mappings from three species, the BERT model was trained on the dataset and Prediction on the reserved positive samples of the three species.b. Comparisons across datasets from three species: The BERT model was trained on Dataset 1.0 and Dataset 0.8, and its performance (ACC) was validated across different datasets and features. c. Performance comparison of different models: Recognition performance (AUC) of BERT, LLAMA3.1-8b, and Qwen2.5-7B on the DNA Dataset 1.0.d, e, f Comparison of predictive probabilities of BERT models trained on Dataset 1.0 and Dataset 0.8 for peak regions: d. Predictive probabilities for 801 peak regions in D. melanogaster.e. Predictive probabilities for 297 peak regions in the first replication of A. thaliana.f. Predictive probabilities for 376 peak regions in the second replication of A. thaliana.
7 The impact of dataset cleaning on the model
To further validate the application potential of the identified methylation motifs, we applied a threshold of 0.80 for motif occurrence in positive samples as the selection criterion (Supplementary Tables 3-5). Using this threshold, we randomly selected negative samples of equal size to the positive samples from various species’ genomes, forming a new dataset, Dataset 0.8, together with the original positive samples. We then trained models on these datasets and compared their performance. As shown in Fig. 4b, compared to the original dataset (Dataset 1.0), the new dataset (Dataset 0.8) demonstrated significant improvements across various evaluation metrics for all three species, whether using the mapped dictionary (164-feature) or the four nucleotide features (A, T, C, G) for training. This indicates that the second dataset, with a motif threshold of P(pos)$\geq 0.80$, was cleaner, leading to improved algorithm performance in identifying methylation sites, as reflected in the ACC. This further demonstrates that the identified motif patterns can capture essential methylation information even without prior knowledge of methylation sites.
8 A Study on Performance Differences in DNA Information Recognition Between Pre-trained and Generative Large Language Models
In the previous section, we verified that the cleaned dataset can be used to train better models. However, in natural language processing, generative large language models with more parameters than the BERT model have now emerged. Therefore, we also compared the performance of these models in recognizing DNA 6mA methylation sites. As shown in Fig. 4c, although the performance of the most advanced generative large language models is high, it still does not surpass the BERT model, which is based on understanding. Therefore, we recommend that the industry select large language models appropriately based on the specific task type.
9 Research on the Practical Application of Models
In practical applications, we further employed the BERT models trained on the two datasets before and after cleaning to predict three datasets derived from D. melanogaster and A. thaliana. These datasets, obtained through DNA immunoprecipitation and next-generation sequencing, represent 6mA-DNA-IP-Seq data, where the exact locations of adenine methylation cannot be determined. We applied the model trained using the 164-feature representation (dict1) to predict methylation sites on the D. melanogaster and A. thaliana 6mA-DNA-IP-Seq data. As shown in Fig. 4d, The BERT trained on dataset 0.8 (the cleaner data) demonstrates significant improvements in predicting methylation sites.
10 Discussion and conclusion
We mapped DNA adenine methylation into the linguistic space of Chinese, enabling the first semantic-level encoding of DNA into a human language. It was found that DNA, like human language, exhibits redundancy. Specifically, DNA possesses an embedded redundant information storage mechanism that helps maintain dynamic stability in liquid media, thereby supporting stable information transfer between DNA and DNA-binding proteins. Moreover, an effective cleaning of the DNA dataset was accomplished. In conclusion, we have proposed a highly innovative approach that addresses the challenges of DNA dataset cleaning and information storage by integrating natural language processing. This approach attempts to understand DNA encoding information from a "linguistic" perspective, even in the absence of known recognition sequences. The discovery of this novel encoding mechanism is expected to offer new insights into biological information systems.
To further synthesize our findings, we constructed a comparative framework highlighting the functional parallels between redundancy in human language and DNA sequences (Supplementary Table 27, Fig.5). In both systems, redundancy serves not merely as excess but as a functional mechanism enhancing error tolerance, adaptive flexibility, and context-dependent disambiguation. For instance, synonymous expressions in language ensure comprehension even under noisy conditions—mirrored by multiple redundant binding motifs in DNA that preserve gene regulation despite mutations. Likewise, linguistic substitutability (e.g., “I’m hungry” vs. “Shall we eat?”) parallels motif diversity in DNA, enabling dynamic regulatory responses under changing environments. These analogies suggest that redundancy is not only structurally embedded but also informationally optimized in both language and genome systems, contributing to stability, adaptability, and semantic nuance.
Fig. 5: Conceptual mapping of redundancy functions between human language and DNA. This model illustrates how three core types of redundancy in natural language—error tolerance, expressive flexibility, and contextual disambiguation—correspond to analogous mechanisms in DNA sequences. These include tolerance to mismatch, multiple motifs supporting the same biological function, and the regulatory role of sequence context. Together, these features contribute to a robust, redundant architecture that supports information stability, signal transmission, and epigenetic memory in biological systems.
When redundancy occurs in a system, it will bring composability to the system, effectively increasing the variety of possible system states. It can be understood that enzymes will have multiple "embrace" forms combined with DNA double-stranded[15, 16], and this different "embrace" situation also implies slight differences in enzyme efficiency. On the contrary, the elimination of redundancy reduces composability and minimizes the number of system states, potentially leading to unintended off-target effects. This redundancy endows DNA with defense, scalability, and adaptability. It enables the system to cope with diverse and unpredictable environmental conditions. This provides the most reasonable explanation for chromatin’s highly dynamic changes and partial destruction during DNA replication, as well as the stability of epigenetic markers that cells are able to maintain throughout the cell cycle[1, 17, 18, 19].
The latest reports suggest that memory formation results from DNA damage followed by repair in response to external environmental stimuli[20]. This process aligns with the discovery of redundant DNA information highlighted in this research, and may represent the essence of epigenetic inheritance and memory[21]. Traits governed by DNA methylation represent a form of inheritance that is susceptible to environmental influences, can be acquired and transmitted across generations, though it may be transient over successive lineages[22, 23]. At present, the mechanisms underlying acquisition and loss remain unclear. Evidence suggests that environment-induced epigenetic mutations and changes are non-random, with the environment driving alterations in specific genomic regions[24]. Jovasevic’s research group found that mice exposed to electric shock (an environmental stressor) activated TLR9 inflammatory signaling, which subsequently induced an immune process within the cell, leading to DNA damage and its repair. Notably, this process did not result in pathogenicity but instead modified the DNA sequence to form memory20. Their analysis further highlighted that unmethylated CpG DNA sequences can activate TLR9 signaling and may also result from DNA demethylation. Jovasevic’s research group research on mice aligns with the general process of epigenetic acquisition: environmental stimuli, DNA modification, sequence alterations, and the formation of long-term memory. Numerous studies have shown that environmental stimuli can induce changes in DNA methylation, resulting in hypermethylation (increased DNA methylation) or hypomethylation (DNA demethylation)[25, 26, 27, 28]. Research indicates that in nondividing cells, DNA can activate DNA demethylation processes to dynamically respond to external stimuli and establish new cellular states[29]. The redundant information features of DNA methylation identified on fixed-length DNA sequences in this study suggest that the interaction between the environment and epigenetics can be interpreted through the dynamics of DNA methylation gain and loss, as illustrated below(Fig. 6). The acquisition of
Fig. 6: Sequence-based information processing system for signal stability and cellular memory. This figure illustrates how environmental stimuli can affect DNA at the sequence level, including structural disruption, insertion or removal of redundant segments, and dynamic changes in epigenetic modifications such as 6mA methylation. Higher redundancy improves signal robustness under fluctuating conditions and supports the long-term maintenance of cellular memory, even in the presence of large-scale chromatin restructuring.(Figure was created with Biorender.com)
epigenetic traits,such as DNA methylation occurs through a series of events: ongoing environmental stimuli trigger immune responses, induce DNA damage and repair, and result in the introduction of new nucleobases. Persistent environmental exposure amplifies DNA methylation signals, leading to redundancy. As redundancy increases, DNA methyltransferases are more likely to methylate the sequence, reinforcing its structural and functional integrity. In turn, higher redundancy leads to more durable memory, supporting the long-term inheritance of adaptive traits. Consequently, the offspring exhibit a consistent genetic predisposition toward a specific personality trait. As for the epigenetic loss of features—characterized by DNA demethylation—this involves the loss of information and reduction of redundancy. Brief environmental stimuli often signify the dominance of another, more persistent environmental factor. In non-dividing cells, this process triggers repeated DNA demethylation, immune reactions, DNA damage, and repair, thereby simplifying the information encoded in DNA. It modifies individual nucleotides, disrupts pre-existing redundancy, reduces methylation signals, and complicates the labeling process by DNA methyltransferases, leading to off-target effects. Consequently, memory data is gradually erased, and traits governed by epigenetic mechanisms weaken across generations, while environmental pressures encode alternative trait information.
The major restriction endonuclease systems, including types I, II, III, and IIG, are always accompanied by methyltransferases. Except for Type II restriction endonucleases (where it is still unclear whether the endonuclease and methyltransferase share the same binding site), the other three types of restriction endonucleases and methyltransferases share the same DNA binding site but cut outside it[30]. Even in the case of Type IV restriction endonucleases, which lack methyltransferase activity, the modifications they rely on are of the methylation type, and their cleavage site lies outside of the binding site, cutting at 12/16 bp[31]. Cleavage within the binding site by restriction endonuclease type II may serve additional functions, such as the rapid elimination of obsolete information or transferring information.
Although the 6mA data from SMRT sources are rigorously filtered and de-duplicated, false-negative samples are still difficult to completely remove. The recognition sites in positive samples are always present at low percentages, which may relate to the DNA information transmission and compensation mechanism identified in this study. When the information is not precise, incomplete, or replaced (Supplementary Tables 3-5), it evidently leads to off-target effects. This means there are always some cases of methylation where low efficiency of the methyltransferases incidentally leads to the methylation of A. It is worth noting that the off-target phenomenon may be nearly impossible to completely avoid on DNA. This means there is always some 6mA where the methylase is in a low-efficiency state, leading to the methylation of adenine.
Information transmission is inevitably subject to noise interference. To ensure accuracy, the sender must encode the information redundantly or repeat it, helping the receiver accurately interpret and complete the transmission. The discovery of redundancy mechanisms in DNA information marks only the beginning of deciphering genomic information. A vast amount of genomic information still needs to be explored and analyzed, which could deepen humanity’s understanding of itself and other life forms. In summary, our findings will provide significant insights and implications for fields such as genomics, synthetic biology, information storage, and gene editing.More importantly, this framework offers a promising direction for developing cross-disciplinary tools that bridge molecular biology, natural language processing, and information theory—opening new avenues in precision medicine, genomic computing, and biological data engineering.Additionally, we observed that some redundant structures may exhibit potential functional directionality in other modification systems, suggesting that they might also play a role at the regulatory level.
Methods
Data Selection and Processing
In biology, there is a process similar to language comprehension: DNA methyltransferases can specifically recognize and methylate methylation sites on DNA. Given the richness of existing adenine methylation data, this study selects DNA6MA methylation data to represent DNA. For species, we chose two multicellular eukaryotes, A. thaliana and D. melanogaster, as well as a prokaryotic organism, Xanthomonas oryzae, the rice pathogen. Research shows that a DNA sequence length of 41 base pairs(bp) for adenine methylation often yields good prediction results[12]. Therefore, in this study, all samples were selected with a sequence length of 41 bp. Any DNA sequence in the dataset can be expressed as:
$S_{\xi}(C)=N_{1}N_{2}\cdots N_{19}N_{20}A_{21}N_{22}N_{23}\cdots N_{+(\xi-1)}N_{+\xi}$
where A is adenine in the center of the DNA sample sequence. When $\xi$ = 41, the experiment achieves the best performance, with each sequence in both positive and negative samples having a total length of 41 and methylated adenine most centrally positioned at location 21.
Positive Sample and Cross-Validation Data Acquisition
Sequences containing the 6mA site for A. thaliana and D. melanogaster were collected from the MethSMRT database[32], whereas data for Xoc BLS256 were obtained from a published study[33]. Optimal predictive accuracy is achieved with segment lengths of 41 bp, positioning the 6mA site centrally; therefore, all positive samples comprised 41 bp sequences. To construct a high-quality dataset, two critical procedures were implemented. Initially, sequences with the identification quality value (Qv) of at least 30 were considered modified sequences, which were retained for further analysis. To reduce redundancy and lessen homology bias, sequences with more than 80% similarity were excluded via the CD-HIT program[34]. These measures ensured the development of precise and objective positive datasets for the stated species. For subsequent cross-validation, we randomly partitioned 20% of the positive samples for future cross-validation, referring to these positive samples as reserved positive samples.
Negative Sample Acquisition from the Genome
Negative samples were randomly collected from the genomes of the three species mentioned above. The selection criteria for negative samples were as follows: first, the adenine at position 21 in these negative samples must not have been experimentally verified as methylated. Second, the negative samples needed to maintain a minimum distance of 20 nucleotides away from any known methylation sites. Finally, the number of negative samples collected was equal to the number of positive samples (excluding the reserved positive samples).
Dataset 1.0
The dataset constructed from the initially collected positive and negative samples is referred to as Dataset 1.0 in this study. Details of the relevant dataset are recorded in Supplementary Table 1.
4 Features description
As the 4 Features description was not changing the DNA sequence, it means that all DNA sequences only have four features(A, T, C, and G).
164 features description
To explore the correlation between DNA and natural language, we hypothesize that fixed-length DNA sequences also exhibit the redundancy characteristic of linguistic symbols. In natural language, Chinese is known for its redundancy, and its square-shaped characters resemble chemical formulas, making Chinese a suitable symbolic system for representing DNA sequences. It has been confirmed that bases at various positions undergo structural changes due to enzyme interactions during methylation[35, 36, 37]. In the methylation reaction process, we consider it inaccurate to represent DNA sequences solely with the four symbols A, T, C, and G. We propose that even the same base should be represented differently based on their positions during the methylation process, much like the saying, "You are not who you were at that time." Therefore, we use Chinese characters to represent the "transitional states" of bases at different positions during methylation.
Before inputting the DNA sequence into BERT model, we first converted the nucleotide sequences into Chinese characters by dictionary. Thus, every DNA sequence contains 164 features (4 * 41 = 164). We prepared three dictionaries with similar functions, corresponding to dict1, dict2, and dict3.
Selection and Training of Algorithm Model
The BERT model[38] is a representative natural language processing understanding model, also recognized as a comprehension-based large language model. Its core idea is to extract contextual information from text through extensive unsupervised learning. This pre-training process endows the model with inherent linguistic meaning. We assume that the information in fixed-length DNA sequences is also redundant, similar to how Chinese has redundancy, and Chinese characters with pictographic features can be likened to chemical formulas. In this study, we chose the Chinese BERT pre-trained model as the deep learning model.
In recent years, BERT has gained prominence as a pre-trained language representation model in natural language processing research. Its key feature is the unsupervised pre-training on large-scale corpora, which allows it to acquire linguistic knowledge embedded in the data. For instance, the Chinese pre-trained BERT model provides a feature vector space for Chinese characters, covering such information as characters, words, semantics, and syntax. Thus, we use thepre-trained Chinese BERT model (for character-level NLP) as a research tool to evaluate the performance of our Dataset 1.0 before and after cleaning by fine-tuning the model, and utilize deep learning to identify potential meaningful information in DNA sequences.
We input data with 4 features and 164 features into the model for training. Since the data in the training set is limited and valuable, we conducted pre-experiments to determine the model’s hyperparameters. We then split each species’ Dataset 1.0 into a training set and a testing set in an 8:2 ratio, ensuring that the training set and testing set do not overlap. We performed 5-fold cross-validation to avoid errors in model training. This process is illustrated in the supplementary Fig.1, with detailed hyperparameter settings as follows:-max_seq_length=50,-train_batch_size=32, -learning_rate=2e-5,-num_train_epochs=8.0. All algorithms were trained and evaluated based on 5-fold cross-validation.
Performance evaluation
We used the following metrics to evaluate the model’s performance: the receiver operating characteristics (ROC) and AUC (the area under ROC), accuracy (ACC), Precision (P), Recall (R), F1, and Matthews correlation coefficient (MCC). The definitions of these metrics are defined as follows:
$P=\frac{TP}{TP+FP}$
$R=\frac{TP}{TP+FN}$
$\frac{1}{F_{1}}=\frac{1}{2}\left(\frac{1}{P}+\frac{1}{R}\right)$
$ACC=\frac{TP+TN}{TP+TN+FN+FP}$
$MCC=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FN)(TP+FP)(TN+FP)(TN+FN)}}$
where TP, TN, FP, and FN represent true positive, true negative, false positive, and false negative, respectively.
Feasibility Verification Experiment of Simulating Base Transition States with Chinese Characters
In this study, we specifically propose that using only the four symbols A, T, C, and G to represent DNA sequences during the methylation process is inaccurate. We hypothesize that DNA, as a language of life, shares redundancy with human language. By mapping DNA into a human language (Chinese) feature space through dictionary mapping, we aim to explore this concept. To verify the feasibility of this approach, we input 4 features (preserving the original information representation) and 164 features into the same model (BERT) for 5-fold cross-validation and evaluation to assess whether any information is lost during the transformation. The results indicate that our proposed approach is feasible, showing that the original symbolic representation may indeed struggle to convey the true information state during the methylation process.
Experiment on the Impact of Different Linguistic Feature Representations on Information
Through the feasibility verification of simulating base transition states with Chinese characters, we found that it is feasible to represent DNA with Chinese characters. Next, we aimed to investigate whether different Chinese character representations would have different effects on the information within DNA sequences. Thus, we used Dictionary 1, Dictionary 2, and Dictionary 3 to map DNA. After performing PCA dimensionality reduction, the three dictionaries are shown in Fig. 1d. Each dictionary-mapped representation was then input as 164 features into the BERT model, followed by 5-fold cross-validation and evaluation. The results further validated our approach, indicating that during the methylation process, each nucleotide undergoes changes in its symbolic representation within the information system.
Information Extraction and Storage Methods on DNA
After training with the same deep learning model, we found that simulating base transition states by converting them into Chinese characters does not destroy the original sequence information. However, the DNA sequences converted to Chinese characters facilitate feature association rule mining. For the association algorithm, we chose Apriori, a classic algorithm in association rule learning, proposed by R. Agrawal et al. in 1994, which is influential in mining frequent itemsets for Boolean association rules[39]. The dataset used was Dataset 1.0 as mentioned above. For the parameters required by Apriori, we set the support to 0.06 and the confidence to 1.0. The mined rules indicate that when certain key nucleotides appear, the central adenine (A) is methylated. Furthermore, these rules must account for at least 6% of the positive samples.
To highlight key information, we calculated the proportion of motifs identified in the positive samples of Dataset 1.0 used for model training. This reflects the tendency of motifs to appear in different categories. The results are presented in Supplementary Tables 3–5, and the calculation formula is as follows:
$P(pos)=\frac{N(pos)}{N(pos)+N(neg)}$ (1)
Where:
• P(pos) represents the proportion of a motif in positive samples.
• N(pos) indicates the number of occurrences of a motif in positive samples.
• N(neg) indicates the number of occurrences of a motif in negative samples.
We also applied a threshold of P(pos)=0.80 to analyze the information storage patterns within individual sequences. These data can be accessed at Supplementary Data. As shown in Figure 2, the storage patterns of sequences with P(pos)$\geq$0.90 within single sequences are displayed. The sequence logos were generated using WebLogo[40].
Introduction of Genomic Background Data
To highlight key information, we computed the recognition rule information (motifs) in the positive and negative samples of Dataset 1.0, effectively incorporating genomic background data to emphasize the information, with the results shown in the supplementary tables. We also used a threshold of 0.9 to determine the form of information storage within individual sequences (as shown in Fig. 2). The threshold of 0.9 indicates that under the influence of genomic background, if a single DNA sequence contains these motifs, the corresponding adenine is likely to be methylated with a certain probability.
Gradient Cleaning Verification Experiment
As mentioned in the introduction, understanding how information is stored on DNA is essential for determining how to clean DNA datasets. The above experiments show that linguistic transformation does not negatively impact the information in DNA sequences, but it greatly facilitates rule mining. Based on rule mining and the introduction of genomic background data, we attempted to address the challenge of cleaning DNA datasets. Using the thresholds (0.95, 0.90, 0.85, 0.80, 0.75) for positive and negative sample ratios from the supplementary tables, we identified false negative samples and used randomly removed equal quantities of negative samples as a control. We performed 5-fold cross-validation for each comparative experiment, with the results shown in Fig. 3a and 3b. The counts of positive and negative samples after removing false negatives are recorded in Fig. 3c.
Cross-Species Validation Experiment
To observe information differences between species, we tested the models trained on Dataset 1.0 using the reserved positive samples from three species and generated a heatmap (Fig. 4a). Based on the information we mined (supplementary tables), we provided a deeper mechanistic explanation.
Comparative Experiment on the Construction of Dataset 0.8 and Model Training with Dataset 1.0
Based on the motifs extracted from the rule mining, we used a threshold of 0.80 for the motifs (Supplementary Tables 3-5) in positive samples as the selection criterion. We then randomly acquired an equal number of negative samples from the genomes of each species. These negative samples do not contain these motifs (P(pos)$\geq$80in positive samples) and are at least 20 nucleotides away from any known methylation sites. Combined with the previous positive samples, this formed a new Dataset 0.8 (Dataset 0.8 in Supplementary Table 2). Like Dataset 1.0, Dataset 0.8 has an equal number of positive and negative samples. We then tested the effect of inputting 4 features and 164 features on the BERT model. Before training, the dataset was split into training and test sets in an 8:2 ratio, followed by 5-fold cross-validation and evaluation. The results are shown in Fig. 4b.
Comparison Experiment Between Models
Thanks to the joint efforts of industry, academia and research, large language models have emerged as a powerful force in artificial intelligence. They have been extensively applied in critical fields such as law, healthcare, and finance, demonstrating powerful cross-domain potential and value. Among them, Llama3.1-8B[41] and Qwen2.5-7B[42] are currently among the top-performing open-source large language models, showcasing strong text processing capabilities in natural language processing, multilingual support, and open-domain question answering.
In our research, we found that inputting 4 features and 164 features does not affect the information within the sequences. Therefore, in the model comparison experiment, all models were standardized with the DNA 4-symbol representation as input to fairly compare different algorithmic models. The models were evaluated on the new Dataset 0.8 using 5-fold cross-validation, with results shown in Fig. 4c.
Model Application
melanogaster and A. thaliana. The 6mA-DNA-IP-Seq data were described in ref.[43] . While 6mA-DNA-IP-Seq does not determine the exact methylation sites of 6mA, it is currently one of the most advanced sequencing techniques with a low error rate. Two replicates with 297 and 376 peaks in A. thaliana and 806 peaks in D. melanogaster were collected. All adenine sites in the 50 bp flanking background around the peaks were selected as prediction objects. The predicted scores of the forward and reverse strands of each 6mA peak with our Bert models were trained on dataset 1.0 and dataset 0.8, respectively. Multiple adenines may be located near a peak, and the maximum score was chosen as the peak score. Their scores were recorded in Supplementary Tables. The model trained on Dataset 0.8 performed much better than the model trained on dataset 1.0(Fig. 4d,e,f), which indicates that we can effectively clean DNA data.
This paper has supplementary information.
Correspondence and material requests should be addressed to db.wang@njau.edu.cn .
Data and Code availability
The data that support the findings of this study are available from GitHub. The GitHub repository will be made publicly available upon formal publication.
Acknowledgments
This work was supported by the Major Project of the National Social Science Fund of China (Project No. 21&ZD331), which enabled the development of this research. We would like to express our sincere gratitude to the funding agency for their trust and support. We also thank all colleagues for their invaluable expertise and assistance throughout the course of this project.